OpenStudy (anonymous):

1 2/3 divided by 3/4

OpenStudy (anonymous):



OpenStudy (anonymous):

6/3

OpenStudy (anonymous):

what lol

OpenStudy (anonymous):

ohhh yea

OpenStudy (johnweldon1993):

\[\large 1 \frac{2}{3} = \frac{3\times 1 + 2}{3} = \frac{5}{3}\] So \[\large \frac{\frac{5}{3}}{\frac{3}{4}} = \frac{5}{3} \times \frac{4}{3} = \frac{20}{9}\]

OpenStudy (anonymous):

Hey @johnweldon1993 Just a tip, when doing an equation like that you can use \div to get a divided by sign. It'd look a little neater than diong a fraction over a fraction So it would look like: \[\frac{5}{3} \div \frac{3}{4}\] Rather than \[\frac{\frac{5}{3}}{\frac{3}{4}}\] Just a little tip.

OpenStudy (johnweldon1993):

No I know, I never liked the look of the division sign however. Thank you though

OpenStudy (anonymous):

Ok. no prob. Just thought I'd let you know.